OpenStudy (anonymous):

Solve by any method. 2x^2+12x=14

OpenStudy (tkhunny):

What method would you like to use? Let's see what you can do.

OpenStudy (anonymous):

Well to get the 14 over, you have to set it up as 2x^2+12-14=0

OpenStudy (tkhunny):

Good call, but you lost an 'x'. Make your life easier and divide by 2. Then what?

OpenStudy (anonymous):

therefore would it be x^2+6x-7=0

OpenStudy (tkhunny):

Perfect. You're almost done. Factoring?

OpenStudy (anonymous):

(x-6)(x+1)?

OpenStudy (tkhunny):

?? "+7" and "-1"

OpenStudy (anonymous):

x? dear lord I hate factoring.

OpenStudy (anonymous):

(x+7)(x-1)= 0 x= -7, 1?

OpenStudy (tkhunny):

Don't blame Him. I think you have it!

OpenStudy (anonymous):

THANK YOU!!!